the time taken by a particle performing shm to pass from pt a to b where its velocities are same is 2s.after another 2 secs it returns to b.the time period of oscillation is?
 
 

Grade:12

2 Answers

Subrata Dutta
8 years ago
According to the question, a and b points are such that they are located at same distances from the equilibrium position. Here also it is said that velocity is same ,i.e; not only the magnitude but also the directions are same. So, total time period of oscillation = 2×( time taken to go from a to b + the next time taken to return at b) = 2×(2+2)= 8 sec(ans) 
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
According to the question, a and b points are such that they are located at same distances from the equilibrium position. Here also it is said that velocity is same ,i.e; not only the magnitude but also the directions are same. So, total time period of oscillation = time taken to go from a to b + the next time taken to return at b + time taken to go from b to a + time taken to return to a = 2 x (time taken to go from a to b + the next time taken to return at b) = 2×(2+2) = 8 sec
 
Hope it helps.
Thanks and regards,
Kushagra